### v1

# On some states minimizing uncertainty relations

Abstract
Analyzing Heisenberg–Robertson (HR) and Schrödinger uncertainty relations we found, that there can exist large set of states of the quantum system under considerations, for which the lower bound of product of the standard deviations of a pair of non–commuting observables, $A$ and $B$, is zero. These states are not eigenstates of either the observable $A$ or $B$. We have also shown that the so–called "sum uncertainty relations" also do not provide any information about lower bounds on the standard deviations calculated for these states.
Introduction
The Heisenberg uncertainty relations [1, 2] was one of milestones in understanding and interpreting the quantum world. In the general, widely accepted case, the quantum uncertainty principle is understood as a lower bound on the product of the standard deviations $\Delta_{\phi}A$ and $\Delta_{\phi}B$ of two non–commuting observables $A$ and $B$ calculated for a given state, say $|\phi\rangle$,
$\Delta_{\phi}A\,\cdot\,\Delta_{\phi}B\geq c>0.$ (1)
Such an interpretation follows from the derivation of the uncertainty relation made by Robertson [3] and Schrödinger [4], (see also [5, 6]). In a general case for an observable $F$ the standard deviation is defined as follows
$\Delta_{\phi}F=\|\delta F|\phi\rangle\|\geq 0,$ (2)
where $\delta F=F-\langle F\rangle_{\phi}\,\mathbb{I}$, and $\langle F\rangle_{\phi}\stackrel{{\scriptstyle\rm def}}{{=}}\langle\phi|F|\phi\rangle$ is the average (or expected) value of an observable $F$ in a system whose state is represented by the normalized vector $|\phi\rangle\in{\cal H}$, provided that $|\langle\phi|F|\phi\rangle|<\infty$. Equivalently: $\Delta_{\phi}F\equiv\sqrt{\langle F^{2}\rangle_{\phi}-\langle F\rangle_{\phi}^{2}}$. The observable $F$ is represented by hermitian operator $F$ acting in a Hilbert space ${\cal H}$ of states $|\phi\rangle$. In general, the relation (1) results from basic assumptions of the quantum theory and from the geometry of Hilbert space [5, 6].
Probably the most common form of inequality (1) is
$\Delta_{\phi}A\cdot\Delta_{\phi}B\;\geq\;\frac{1}{2}\left|\langle[A,B]\rangle_{\phi}\right|,$ (3)
which holds for any two observables, $A$ and $B$, represented by non–commuting hermitian operators $A$ and $B$ acting in the Hilbert space of states (see [3] and also [4, 5, 6]), such that $[A,B]$ exists and $|\phi\rangle\in{\cal D}(AB)\bigcap{\cal D}(BA)$, (${\cal D}({\cal O})$ denotes the domain of an operator $\cal O$ or of a product of operators). The derivation of inequality (3) is the rigorous one. Indeed, the first step is to use the Schwartz inequality
$\left\|\;|\psi_{A}\rangle\right\|\,\left\|\;|\psi_{B}\rangle\right\|\geq\left|\langle\psi_{A}|\psi_{B}\rangle\right|,$ (4)
and write it appropriately for vectors $|\psi_{A}\rangle=\delta A|\phi\rangle$ and $|\psi_{B}\rangle=\delta B|\phi\rangle$ (see, e.g. [6]):
$\left\|\delta A|\phi\rangle\right\|^{2}\;\left\|\delta B|\phi\rangle\right\|^{2}\,\geq\,\left|\langle\phi|\delta A\;\delta B|\phi\rangle\right|^{2},$ (5)
which holds for all $|\phi\rangle\in{\cal D}(AB)\bigcap{\cal D}(BA)$. (The equality in (4), or in (5), holds when $|\psi_{A}\rangle=z|\psi_{B}\rangle$, $z\in\mathbb{C}$). The next step is to transform the right side of Eq. (5) as follows:
$\left|\langle\phi|\delta A\;\delta B|\phi\rangle\right|^{2}$ $=$ $\left[\Re\,(\langle\phi|\delta A\;\delta B|\phi\rangle)\right]^{2}+\left[\Im\,(\langle\phi|\delta A\;\delta B|\phi\rangle)\right]^{2},$ (6)
$=$ $\frac{1}{4}\left(\langle\phi|(\delta A\;\delta B\,+\,\delta B\;\delta A)|\phi\rangle\right)^{2}$
$\;\;\;\;\;\;\;+\;\frac{1}{4}\left|\langle\phi|(\delta A\;\delta B\,-\,\delta B\;\delta A)|\phi\rangle\right|^{2}$
$\equiv$ $\frac{1}{4}\left(\langle\phi|(\delta A\;\delta B\,+\,\delta B\;\delta A)|\phi\rangle\right)^{2}\,+\,\frac{1}{4}\left|\langle\phi|[A,B]|\phi\rangle\right|^{2}$
$\geq$ $\frac{1}{4}\left|\langle\phi|[A,B]|\phi\rangle\right|^{2},$ (7)
where $\Re\,(z)$ denotes the real part of the complex number $z$ and $\Im\,(z)$ is the imaginary part of $z$. The property $[\delta A,\delta B]=[A,B]$ taking place for all $|\phi\rangle\in{\cal D}(AB)\bigcap{\cal D}(BA)$ was used in (6).
Note that if to replace the right hand side of Eq. (5) by (6) then one obtains the uncertainty relation of the type derived by Schrödinger [4]:
$(\Delta_{\phi}A)^{2}\,\cdot\,(\Delta_{\phi}B)^{2}\,\geq\frac{1}{4}\left(\langle\phi|(\delta A\;\delta B\,+\,\delta B\;\delta A)|\phi\rangle\right)^{2}\,+\,\frac{1}{4}\left|\langle\phi|[A,B]|\phi\rangle\right|^{2},$ (8)
or, equivalently, in more familiar form,
$(\Delta_{\phi}A)^{2}\,\cdot\,(\Delta_{\phi}B)^{2}\,\geq\left(\frac{\langle(AB+BA)\rangle_{\phi}}{2}-\langle A\rangle_{\phi}\,\langle B\rangle_{\phi}\right)^{2}\,+\,\left|\frac{\langle[A,B]\rangle_{\phi}}{2}\right|^{2}.$ (9)
As it can be seen relations (8), (9) seem to be more general and precise than the relation (3). It is because the right hand sides of (8), (9) are strictly equivalent to the right hand side of Eq. (5).
Now if one replaces the right hand side in Eq. (5) by (7) then one obtains the uncertainty relation (3) as a result. The transformations leading to formulas (7), (6) prove that the Heisenberg–Robertson (HR) uncertainty relation (3) is less precise than the Schrödinger uncertainty relations (8), (9).
Results
Let us analyze now properties of inequalities (3) and (5) — (9). Firstly, as is well known, the right side of (3) becomes zero when $[A,B]\neq 0$ and vector $|\phi\rangle$ is an eigenvector of $A$ or $B$. If to assume that, e.g., $|\phi\rangle=|\psi_{b}\rangle$ is a normalized eigenvector of $B$ for the eigenvalue $b$, that is that $B|\psi_{b}\rangle=b|\psi_{b}\rangle$, then one immediately finds that $\langle\psi_{b}|B|\psi_{b}\rangle=b$, and $\langle\psi_{b}|AB|\psi_{b}\rangle=b\langle\psi_{b}|A|\psi_{b}\rangle$ and also $\delta_{\psi_{b}}B|\psi_{b}\rangle\equiv 0$. As a result we have that $\langle\psi_{b}|[A,B]|\psi_{b}\rangle\equiv 0$ and that $\Delta_{\psi_{b}}(B)=0$. The same effect, i. e. that the right hand side of (3) is equal to zero for $|\phi\rangle$ being an eigenvector of $A$ or $B$, takes place in the case of more general inequalities (8), (9). It is because, the two sides of the Schwartz inequality (4) are equal to zero if $|\psi_{A}\rangle=0$ (or $|\psi_{B}\rangle=0$), and in the considered case $|\phi\rangle=|\psi_{b}\rangle$ there is $|\psi_{B}\rangle\equiv\delta_{\psi_{b}}B|\psi_{b}\rangle\equiv 0$ in (5).
The second, much more interesting, non–trivial case of the vanishing right side of the Schwartz inequality (4), and thus the inequality (5), is when $[A,B]\neq 0$ and vectors both appearing therein $|\psi_{A}\rangle=\delta A\,|\phi\rangle,|\psi_{B}\rangle=\delta B\,|\phi\rangle$ are nonzero and $\delta A\,|\phi\rangle\,\perp\,\delta B\,|\phi\rangle$. This implies that there are $\Delta_{\phi}A>0,\,\Delta_{\phi}B>0$ and $\Delta_{\phi}A\,\cdot\,\Delta_{\phi}B\geq|\langle\phi|\delta A\,\delta B|\phi\rangle|\equiv 0$ instead of inequalities (1), (3) and (8), (9). So in this case there are no lower bounds for product of standard deviations $\Delta_{\phi}A$ and $\Delta_{\phi}B$. More precisely: they are equal to zero. The number of such $|\psi_{A}\rangle=\delta A\,|\phi\rangle,|\psi_{B}\rangle=\delta B\,|\phi\rangle$, strictly speaking the number of corresponding vectors $|\phi\rangle$, can be extremely large depending on the dimension of the state space. As it can be easily see there must be: $\langle\phi|\psi_{A}\rangle=\langle\phi|\delta A\,|\phi\rangle\equiv 0$ and similarly $\langle\phi|\psi_{B}\rangle=\langle\phi|\delta B\,|\phi\rangle\equiv 0$. In other words there are: $|\psi_{A}\rangle=\delta A\,|\phi\rangle\equiv|\psi_{A}^{\perp}\rangle\,\perp\,|\phi\rangle$ and $|\psi_{B}\rangle=\delta B\,|\phi\rangle\equiv|\psi_{B}^{\perp}\rangle\,\perp\,|\phi\rangle$. As we have already seen, we analyze the condition $\langle\psi_{A}|\psi_{B}\rangle\equiv\langle\psi_{A}^{\perp}|\psi_{B}^{\perp}\rangle=0$, that is $|\psi_{A}\rangle\equiv|\psi_{A}^{\perp}\rangle\,\perp\,|\psi_{B}\rangle\equiv|\psi_{B}^{\perp}\rangle$. Summing up: For a given $|\phi\rangle$ we have $|\phi\rangle\perp|\psi_{A}^{\perp}\rangle,\;|\phi\rangle\perp|\psi_{B}^{\perp}\rangle$ and $|\psi_{A}^{\perp}\rangle\perp|\psi_{B}^{\perp}\rangle$. This problem has no non–zero solutions in two–dimensional space. They exist when the dimension of the state space is not less than 3.
As an illustration of the case under consideration, let us analyze the following example: Assume that $A=\lambda_{3}$ and $B=\lambda_{4}$, where $\lambda_{3},\lambda_{4}$ are Gell–Mann matrices:
$\lambda_{3}=\left(\begin{array}[]{ccc}1&0&0\\ 0&-1&0\\ 0&0&0\\ \end{array}\right),\;\;\lambda_{4}=\left(\begin{array}[]{ccc}0&0&1\\ 0&0&0\\ 1&0&0\\ \end{array}\right),\;\;\lambda_{5}=\left(\begin{array}[]{ccc}0&0&i\\ 0&0&0\\ -i&0&0\\ \end{array}\right).$ (10)
Here $\lambda_{5}$ is also Gell–Mann matrix. They are self–adjoint and do not commute, $[\lambda_{3},\lambda_{4}]=-i\lambda_{5}\neq 0.$ For these matrices the inequality (5) takes the following form,
$\left\|\delta\lambda_{3}\,|\phi\rangle\right\|^{2}\,\cdot\,\left\|\delta\lambda_{4}|\phi\rangle\right\|^{2}\geq\left|\langle\phi|\delta\lambda_{3}\,\delta\lambda_{4}|\phi\rangle\right|^{2}.$ (11)
Let us choose for simplicity,
$|\phi\rangle=|\phi_{1}\rangle=\frac{1}{N}\,\left(\begin{array}[]{c}a\\ b\\ 0\\ \end{array}\right),$ (12)
where $N^{2}=|a|^{2}+|b|^{2}$ and $a,b\in\mathbb{C}$. Using (12) one gets
$\delta\lambda_{3}|\phi_{1}\rangle=\frac{1}{N^{3}}\,\left(\begin{array}[]{c}2|b|^{2}a\\ -2|a|^{2}b\\ 0\\ \end{array}\right)\neq 0,$ (13)
and
$\delta\lambda_{4}|\phi_{1}\rangle=\frac{1}{N}\,\left(\begin{array}[]{c}0\\ 0\\ a\\ \end{array}\right)\neq 0,$ (14)
(where $\delta\lambda_{k}|\phi_{1}\rangle=(\lambda_{k}-\langle\phi_{1}|\lambda_{k}|\phi_{1}\rangle)|\phi_{1}\rangle$ and $k=3,4$), which leads to the result $\langle\phi_{1}|\delta\lambda_{3}\,\delta\lambda_{4}|\phi_{1}\rangle=0$, and $(\Delta_{{\phi}_{1}}\lambda_{3})^{2}={\left\|\delta\lambda_{3}|\phi_{1}\rangle\right\|}^{2}\neq 0$, $(\Delta_{{\phi}_{1}}\lambda_{4})^{2}={\left\|\delta\lambda_{4}|\phi_{1}\rangle\right\|}^{2}\neq 0$. Hence one concludes that for $|\phi_{1}\rangle$ the inequality (11) and thus relations (3), (8), (9) take the following form,
$\Delta_{\phi_{1}}\lambda_{3}\,\cdot\,\Delta_{\phi_{1}}\lambda_{4}\geq 0.$ (15)
Note that this result holds for any $a\neq 0$ and $b\neq 0$ defining the vector $|\phi_{1}\rangle$. This shows that the number of such vectors may be very large.
Discussion
Recently, uncertainty relations for sums of standard deviations or variances have been studied in the literature (see, e. g. (see, e. g. [7, 11, 9, 12, 13] and many other papers), which are considered to be more useful than the HR or Schrödinger uncertainty relations. These studies are motivated by the fact that the latter do not provide any bounds on the standard deviations of either the observables $A$ or $B$ when the system is in the state described by the eigenfunction of the observable $B$ (or $A$) which causes that the right-hand side of (3) inequality vanishes. Relations of this type are called "sum uncertainty relations". Let us now look at some of these "sum uncertainty relations" and see if they are sensitive to the effects described above. Their mathematical basis is the triangle inequality in the Hilbert state space: There is
$\left\|\,|\psi_{1}\rangle\right\|+\left\|\,|\psi_{2}\rangle\right\|\,\geq\,\left\|\,|\psi_{1}\rangle+|\psi_{2}\rangle\right\|,$ (16)
for each pair of vectors $|\psi_{1}\rangle,|\psi_{2}\rangle\in\cal{H}$. It is usually derived in Hilbert space using the Schwartz inequality (4) — see e. g. see, e. g., [6, 7, 8, 9]. If we now substitute $|\psi_{A}\rangle=\delta A|\phi\rangle$ and $|\psi_{B}\rangle=\delta B|\phi\rangle$ into (16) instead of $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, we get the "sum uncertainty relation" derived in [7],
$\left\|\delta A|\phi\rangle\right\|+\left\|\delta B\phi\rangle\right\|\,\geq\,\left\|\delta A|\phi\rangle+\delta B|\phi\rangle\right\|\equiv\left\|\delta(A+B)|\phi\rangle\right\|,$ (17)
that is,
$\Delta_{\phi}A+\Delta_{\phi}B\geq\Delta_{\phi}(A+B).$ (18)
When looking for inequalities of the "sum uncertainty relation" type, one can also use the triangle inequality of the second kind [10, 11]:
$\left\|\,|\psi_{1}\rangle\right\|^{2}+\left\|\,|\psi_{2}\rangle\right\|^{2}\,\geq\,\frac{1}{2}\left\|\,|\psi_{1}\rangle+|\psi_{2}\rangle\right\|^{2},$ (19)
which leads to the following sum uncertainty relation,
$(\Delta_{\phi}A)^{2}+(\Delta_{\phi}B)^{2}\geq\frac{1}{2}\left[\,\Delta_{\phi}(A+B)\right]^{2}.$ (20)
This type of generalization of the inequality (18) has been studied in many papers (see, e. g. [11, 12, 13, 14, 15, 16] and many others).
Now, let $|\phi\rangle=|\psi_{b}\rangle$, where $|\psi_{b}\rangle$ is and eigenvector of the observable $B$ for the eigenvalue $b$. Then, as it has been shown earlier, $\delta B|\psi_{b}\rangle=0$, and therefore $\Delta_{\psi_{b}}B=0$, and $\delta(A+B)|\psi_{b}\rangle=\delta A|\psi_{b}\rangle+\delta B|\psi_{b}\rangle\equiv\delta A|\psi_{b}\rangle$, and as a result we have $\Delta_{\psi_{b}}(A+B)\equiv\Delta_{\psi_{b}}A$. This means that even though the right–hand side of the inequalities (18) and (20) is non–zero, these inequalities become trivial: $\Delta_{\psi_{b}}A\geq\Delta_{\psi_{b}}A$ (or $(\Delta_{\psi_{b}}A)^{2}\geq\frac{1}{2}(\Delta_{\psi_{b}}A)^{2}$ respectively), and contains no information about $\Delta_{\psi_{b}}A$.
Let us now examine a case that is more interesting to us: Assume that $|\psi_{A}\rangle=\delta A|\phi\rangle\neq 0,\;|\psi_{B}\rangle=\delta B|\phi\rangle\neq 0$, and $\delta A|\phi\rangle\,\perp\,\delta B|\phi\rangle$. Analyzing the right hand sides of inequalities (19) and (20) with this assumption, one finds that $\left[\Delta_{\phi}(A+B)\right]^{2}\equiv\left\|\delta A|\phi\rangle+\delta B|\phi\rangle\right\|^{2}\equiv\left\|\delta A|\phi\rangle\right\|^{2}+\left\|\delta B\phi\rangle\right\|^{2}$. (It is because $\left\|\,|\psi_{A}\rangle+|\psi_{B}\rangle\right\|^{2}=\left\|\,|\psi_{A}\rangle\right\|^{2}+\left\|\,|\psi_{B}\rangle\right\|^{2}$ for any $|\psi_{A}\rangle\,\perp\,|\psi_{B}\rangle$). This means that finally, in the case under consideration, the inequality (20) takes the following form: $(\Delta_{\phi}A)^{2}+(\Delta_{\phi}B)^{2}\geq\frac{1}{2}\left[(\Delta_{\phi}A)^{2}+(\Delta_{\phi}B)^{2}\right]$ and we have no information about lower bounds for $(\Delta_{\phi}A)^{2}$ and $(\Delta_{\phi}B)^{2}$. Let us now analyze the inequality (18) using the assumption that $\delta A|\phi\rangle\,\perp\,\delta B|\phi\rangle$. We already know that then $\left[\Delta_{\phi}(A+B)\right]^{2}\equiv(\Delta_{\phi}A)^{2}+(\Delta_{\phi}B)^{2}$. Next, by multiplying the two sides of the inequality (18) by each other respectively, we get that $(\Delta_{\phi}A)^{2}+(\Delta_{\phi}B)^{2}+2\Delta_{\phi}A\,\Delta_{\phi}B\geq\left[\Delta_{\phi}(A+B)\right]^{2}\equiv(\Delta_{\phi}A)^{2}+(\Delta_{\phi}B)^{2}$. This inequality simplifies to $\Delta_{\phi}A\,\Delta_{\phi}B\geq 0$, which again gives us no information about the lower bounds on $\Delta_{\phi}A$ and $\Delta_{\phi}B$.
In summary, when we study two non–commuting observables, $A$ and $B$, and apply the "sum uncertainty relations" to find lower bounds on $\Delta_{\phi}A$ and $\Delta_{\phi}B$ in the situations considered above, the conclusions resulting from these relations are the same as the conclusions resulting from the HR uncertainty relations. The "sum uncertainty relations" seem to be more useful in the case where we investigate possible relations between lower bounds of the standard deviations, $(\Delta_{\phi}A_{1}),(\Delta_{\phi}A_{2}),\ldots,(\Delta_{\phi}A_{n})$ of a set of non–commuting observables $\left\{A_{j}\right\}_{j=1}^{n}$ with $n\geq 3$. In such cases, using the generalization of the triangle inequality to $n$ vectors (see, e. g., [8]),
$\sum_{j=1}^{n}\left\|\,|\psi_{j}\rangle\right\|\;\geq\;\left\|\sum_{j=1}^{n}|\psi_{j}\rangle\right\|,$ (21)
we obtain the following generalization of the inequality (17),
$\sum_{j=1}^{n}(\Delta_{\phi}A_{j})\geq\,\Delta_{\phi}(\sum_{j=1}^{n}A_{j}).$ (22)
and this type of "sum uncertainty relation" and its various generalizations analogous to (20) have recently been intensively studied in many papers (see, e. g. [7, 14, 15, 16] and many others).
Final remarks
As it has been shown there may exist large sets of states $\{|\phi\rangle\}_{A,B}\subset{\cal H}$, (where $\dim{\cal H}\geq 3$), of a quantum system, which are not eigenstates of any observable from non–commuting pairs $A$ and $B$, such that vectors $\delta A|\phi\rangle\neq 0$ and $\delta B|\phi\rangle\neq 0$ are orthogonal: $\delta A|\phi\rangle\,\perp\,\delta B|\phi\rangle$. These sets may be different for different pairs of such observables. On these sets, the right–hand sides of the inequalities (3), (5) — (9) reach their absolute minimum, i.e. they have the value zero. So, there is no lower bound for the product $\Delta_{\phi}A\cdot\Delta_{\phi}B$. Strictly speaking, this lower bound is zero. In other words, quantum theory (to be more precisely: basic assumptions (postulates) of the quantum theory ) allows $\Delta_{\phi}A$ and $\Delta_{\phi}B$ both to be as small as possible and also it allows that in this case the magnitude of $\Delta_{\phi}A$ does not affect the magnitude of $\Delta_{\phi}B$ and vice versa. This raises the questions: Can this property be used technically in any way, and if so, how? What are properties of the system prepared in such a state?