### v2

# On some states minimizing uncertainty relations

Abstract
Analyzing Heisenberg–Robertson (HR) and Schrödinger uncertainty relations we found, that there can exist large set of states of the quantum system under considerations, for which the lower bound of the product of the standard deviations of a pair of non–commuting observables, $A$ and $B$, is zero. These states are not eigenstates of either the observable $A$ or $B$. We have also shown that the so–called "sum uncertainty relations" also do not provide any information about lower bounds on the standard deviations calculated for these states.
Introduction
The Heisenberg uncertainty relations [1, 2] was one of milestones in understanding and interpreting the quantum world. In the general, widely accepted case, the quantum uncertainty principle is understood as a lower bound on the product of the standard deviations $\Delta_{\phi}A$ and $\Delta_{\phi}B$ of two non–commuting observables $A$ and $B$ calculated for a given state, say $|\phi\rangle$,
$\Delta_{\phi}A\,\cdot\,\Delta_{\phi}B\geq c>0.$ (1)
Such an interpretation follows from the derivation of the uncertainty relation made by Robertson [3] and Schrödinger [4], (see also [5, 6]). In a general case for an observable $F$ the standard deviation is defined as follows
$\Delta_{\phi}F=\|\delta_{\phi}F|\phi\rangle\|\geq 0,$ (2)
where $\delta_{\phi}F=F-\langle F\rangle_{\phi}\,\mathbb{I}$, and $\langle F\rangle_{\phi}\stackrel{{\scriptstyle\rm def}}{{=}}\langle\phi|F|\phi\rangle$ is the average (or expected) value of an observable $F$ in a system whose state is represented by the normalized vector $|\phi\rangle\in{\cal H}$, provided that $|\langle\phi|F|\phi\rangle|<\infty$. Equivalently: $\Delta_{\phi}F\equiv\sqrt{\langle F^{2}\rangle_{\phi}-\langle F\rangle_{\phi}^{2}}$. The observable $F$ is represented by hermitian operator $F$ acting in a Hilbert space ${\cal H}$ of states $|\phi\rangle$. In general, the relation (1) results from basic assumptions of the quantum theory and from the geometry of Hilbert space [5, 6].
Probably the most common form of inequality (1) is
$\Delta_{\phi}A\cdot\Delta_{\phi}B\;\geq\;\frac{1}{2}\left|\langle[A,B]\rangle_{\phi}\right|,$ (3)
which holds for any two observables, $A$ and $B$, represented by non–commuting hermitian operators $A$ and $B$ acting in the Hilbert space of states (see [3] and also [4, 5, 6]), such that $[A,B]$ exists and $|\phi\rangle\in{\cal D}(AB)\bigcap{\cal D}(BA)$, (${\cal D}({\cal O})$ denotes the domain of an operator $\cal O$ or of a product of operators). The derivation of inequality (3) is the rigorous one. Indeed, the first step is to use the Schwartz inequality
$\left\|\;|\psi_{A}\rangle\right\|\,\left\|\;|\psi_{B}\rangle\right\|\geq\left|\langle\psi_{A}|\psi_{B}\rangle\right|,$ (4)
and write it appropriately for vectors $|\psi_{A}\rangle=\delta_{\phi}A|\phi\rangle$ and $|\psi_{B}\rangle=\delta_{\phi}B|\phi\rangle$ (see, e.g. [6]):
$\left\|\delta_{\phi}A|\phi\rangle\right\|^{2}\;\left\|\delta_{\phi}B|\phi\rangle\right\|^{2}\,\geq\,\left|\langle\phi|\delta_{\phi}A\;\delta_{\phi}B|\phi\rangle\right|^{2},$ (5)
which holds for all $|\phi\rangle\in{\cal D}(AB)\bigcap{\cal D}(BA)$. (The equality in (4), or in (5), holds when $|\psi_{A}\rangle=z|\psi_{B}\rangle$, $z\in\mathbb{C}$). The next step is to transform the right side of Eq. (5) as follows:
$\left|\langle\phi|\delta_{\phi}A\;\delta_{\phi}B|\phi\rangle\right|^{2}$ $=$ $\left[\Re\,(\langle\phi|\delta_{\phi}A\;\delta_{\phi}B|\phi\rangle)\right]^{2}+\left[\Im\,(\langle\phi|\delta_{\phi}A\;\delta_{\phi}B|\phi\rangle)\right]^{2},$
$=$ $\frac{1}{4}\left(\langle\phi|(\delta_{\phi}A\;\delta_{\phi}B\,+\,\delta_{\phi}B\;\delta_{\phi}A)|\phi\rangle\right)^{2}$
$\;\;\;\;\;\;\;+\;\frac{1}{4}\left|\langle\phi|(\delta_{\phi}A\;\delta_{\phi}B\,-\,\delta_{\phi}B\;\delta_{\phi}A)|\phi\rangle\right|^{2}$
$\equiv$ $\frac{1}{4}\left(\langle\phi|(\delta_{\phi}A\;\delta_{\phi}B\,+\,\delta_{\phi}B\;\delta_{\phi}A)|\phi\rangle\right)^{2}\,+\,\frac{1}{4}\left|\langle\phi|[A,B]|\phi\rangle\right|^{2}$ (7)
$\geq$ $\frac{1}{4}\left|\langle\phi|[A,B]|\phi\rangle\right|^{2},$ (8)
where $\Re\,(z)$ denotes the real part of the complex number $z$ and $\Im\,(z)$ is the imaginary part of $z$. The property $[\delta_{\phi}A,\delta_{\phi}B]=[A,B]$ taking place for all $|\phi\rangle\in{\cal D}(AB)\bigcap{\cal D}(BA)$ was used in (7).
Note that if to use definition (2) and then replace the right hand side of Eq. (5) by (7) then one obtains the uncertainty relation of the type derived by Schrödinger [4]:
$(\Delta_{\phi}A)^{2}\,\cdot\,(\Delta_{\phi}B)^{2}\,\geq\frac{1}{4}\left(\langle\phi|(\delta_{\phi}A\;\delta_{\phi}B\,+\,\delta_{\phi}B\;\delta_{\phi}A)|\phi\rangle\right)^{2}\,+\,\frac{1}{4}\left|\langle\phi|[A,B]|\phi\rangle\right|^{2},$ (9)
or, equivalently, in more familiar form,
$(\Delta_{\phi}A)^{2}\,\cdot\,(\Delta_{\phi}B)^{2}\,\geq\left(\frac{\langle(AB+BA)\rangle_{\phi}}{2}-\langle A\rangle_{\phi}\,\langle B\rangle_{\phi}\right)^{2}\,+\,\left|\frac{\langle[A,B]\rangle_{\phi}}{2}\right|^{2}.$ (10)
As it can be seen relations (9), (10) seem to be more general and precise than the relation (3). It is because the right hand sides of (9), (10) are strictly equivalent to the right hand side of Eq. (5).
Now if one replaces the right hand side in Eq. (5) by (8) then one obtains the uncertainty relation (3) as a result. The transformations leading to formulas (8), (7) prove that the Heisenberg–Robertson (HR) uncertainty relation (3) is less precise than the Schrödinger uncertainty relations (9), (10). The uncertainty relations (3) and (9), (10) are state–depended and only these types of relations will be considered in the following parts of the paper.
Results
Let us analyze now properties of inequalities (3) and (5) — (10). Firstly, as is well known, the right side of (3) becomes zero when $[A,B]\neq 0$ and vector $|\phi\rangle$ is an eigenvector of $A$ or $B$. If to assume that, e.g., $|\phi\rangle=|\psi_{b}\rangle$ is a normalized eigenvector of $B$ for the eigenvalue $b$, that is that $B|\psi_{b}\rangle=b|\psi_{b}\rangle$, then one immediately finds that $\langle\psi_{b}|B|\psi_{b}\rangle=b$, and $\langle\psi_{b}|AB|\psi_{b}\rangle=b\langle\psi_{b}|A|\psi_{b}\rangle$ and also $\delta_{\psi_{b}}B|\psi_{b}\rangle\equiv 0$. As a result we have that $\langle\psi_{b}|[A,B]|\psi_{b}\rangle\equiv 0$ and that $\Delta_{\psi_{b}}(B)=0$. The same effect, i. e. that the right hand side of (3) is equal to zero for $|\phi\rangle$ being an eigenvector of $A$ or of $B$, takes place in the case of more general inequalities (9), (10). It is because, the two sides of the Schwartz inequality (4) are equal to zero if $|\psi_{A}\rangle=0$ (or $|\psi_{B}\rangle=0$), and in the considered case $|\phi\rangle=|\psi_{b}\rangle$ there is $|\psi_{B}\rangle\equiv\delta_{\psi_{b}}B|\psi_{b}\rangle\equiv 0$ in (5).
The second, much more interesting, non–trivial case of the vanishing right side of the Schwartz inequality (4), and thus the inequality (5), is when $[A,B]\neq 0$ and vectors both appearing therein $|\psi_{A}\rangle=\delta_{\phi}A\,|\phi\rangle,|\psi_{B}\rangle=\delta_{\phi}B\,|\phi\rangle$ are nonzero and $\delta_{\phi}A\,|\phi\rangle\,\perp\,\delta_{\phi}B\,|\phi\rangle$. This implies that there are $\Delta_{\phi}A>0,\,\Delta_{\phi}B>0$ and $\Delta_{\phi}A\,\cdot\,\Delta_{\phi}B\geq|\langle\phi|\delta_{\phi}A\,\delta_{\phi}B|\phi\rangle|\equiv 0$ instead of inequalities (1), (3) and (9), (10). So in this case there is no a lower bound for product of standard deviations $\Delta_{\phi}A$ and $\Delta_{\phi}B$. More precisely: it is equal to zero. The number of such $|\psi_{A}\rangle=\delta_{\phi}A\,|\phi\rangle,|\psi_{B}\rangle=\delta_{\phi}B\,|\phi\rangle$, strictly speaking the number of corresponding vectors $|\phi\rangle$, can be extremely large depending on the dimension of the state space. As it can be easily see there must be: $\langle\phi|\psi_{A}\rangle=\langle\phi|\delta_{\phi}A\,|\phi\rangle\equiv 0$ and similarly $\langle\phi|\psi_{B}\rangle=\langle\phi|\delta_{\phi}B\,|\phi\rangle\equiv 0$. In other words there are: $|\psi_{A}\rangle=\delta_{\phi}A\,|\phi\rangle\stackrel{{\scriptstyle\rm def}}{{=}}\alpha_{A}|\phi_{A}^{\perp}\rangle\,\perp\,|\phi\rangle$ and $|\psi_{B}\rangle=\delta_{\phi}B\,|\phi\rangle\stackrel{{\scriptstyle\rm def}}{{=}}\alpha_{B}|\phi_{B}^{\perp}\rangle\,\perp\,|\phi\rangle$, where $\|\,|\phi_{A}^{\perp}\rangle\|=\|\,|\phi_{B}^{\perp}\rangle\|=1$, (see also [7]). As we have already seen, we analyze the condition $\langle\psi_{A}|\psi_{B}\rangle\equiv\langle\phi_{A}^{\perp}|\phi_{B}^{\perp}\rangle=0$, that is $|\psi_{A}\rangle\equiv\alpha_{A}|\phi_{A}^{\perp}\rangle\,\perp\,|\psi_{B}\rangle\equiv\alpha_{B}|\phi_{B}^{\perp}\rangle$. Summing up: For a given $|\phi\rangle$ we have $|\phi\rangle\perp|\phi_{A}^{\perp}\rangle,\;|\phi\rangle\perp|\phi_{B}^{\perp}\rangle$ and $|\phi_{A}^{\perp}\rangle\perp|\phi_{B}^{\perp}\rangle$. This problem has no non–zero solutions in two–dimensional space. They exist when the dimension of the state space is not less than 3.
States $|\psi_{A}\rangle=\delta_{\phi}A\,|\phi\rangle,|\psi_{B}\rangle=\delta_{\phi}B\,|\phi\rangle$ have other interesting properties. Analyzing the left side of equation (Introduction) we find
$\langle\phi|\delta_{\phi}A\,\delta_{\phi}B|\phi\rangle\equiv\langle\phi|AB|\phi\rangle-\langle A\rangle_{\phi}\,\langle B\rangle_{\phi}\stackrel{{\scriptstyle\rm def}}{{=}}{\cal C}_{\phi}(A,B),$ (11)
where ${\cal C}_{\phi}(A,B)$ is a correlation function (see, eg. [8, 9]). This means that
$\langle\phi|\delta_{\phi}A\,\delta_{\phi}B|\phi\rangle=0\;\Leftrightarrow\;\langle\phi|AB|\phi\rangle-\langle A\rangle_{\phi}\,\langle B\rangle_{\phi}=0.$ (12)
Now, if we use definitions (2) in (5) and (12) then we get the following inequality equivalent to (9) and (10)
$\Delta_{\phi}A\cdot\Delta_{\phi}B\;\geq\;\left|{\cal C}{\phi}(A,B)\right|,$ (13)
which seems to be the most general uncertainty relation. If the system is in a state $|\phi\rangle$ such that $\Delta_{\phi}A>0$ and $\Delta_{\phi}B>0$, then the inequality (13) can be written in an equivalent, useful in some applications, form as
$\mathfrak{r}_{\phi}(A,B)\stackrel{{\scriptstyle\rm def}}{{=}}\frac{\left|{\cal C}_{\phi}(A,B)\right|}{\Delta_{\phi}A\cdot\Delta_{\phi}B}\leq 1,$ (14)
which looks like a variant of the Pearson correlation coefficient (see, e.g. [9]) or rather its quantum modification. It is easy to find some formal properties of ${\cal C}_{\phi}(A,B)$ and of $\mathfrak{r}_{\phi}(A,B)$:
${\cal C}_{\phi}(A,A)$ $=$ $(\Delta_{\phi}A)^{2},$
${\cal C}_{\phi}(A,B)$ $=$ $[{\cal C}_{\phi}(B,A)]^{\ast},$
${\cal C}_{\phi}(A,B_{1}+B_{2})$ $=$ ${\cal C}_{\phi}(A,B_{1})+{\cal C}_{\phi}(A,B_{2}),$ (15)
$\mathfrak{r}_{\phi}(A,A)=1,\;\;\mathfrak{r}_{\phi}(A,B)=\mathfrak{r}_{\phi}(B,A).$ (16)
It should be noted that $\mathfrak{r}_{\phi}(A,A)=1$ even though $[A,A]$=0, which means that the right side of the HR relation (3) is equal to zero. Moreover, $\mathfrak{r}_{\phi}(A,B)=0$ if only there exists such a state $|\phi\rangle$ that $\delta_{\phi}A|\phi\rangle\perp\delta_{\phi}B|\phi\rangle$ despite the fact that $[A,B]\neq 0$. The result $\mathfrak{r}_{\phi}(A,B)=0$ means that observables $A$ and $B$ are uncorrelated in the state $|\phi\rangle$. However, this does not exclude that in a state $|\psi\rangle\neq|\phi\rangle$ there may be $\mathfrak{r}_{\psi}(A,B)>0$ and then the value $\mathfrak{r}_{\psi}(A,B)$ describes the level of a correlation.
Another observation concerns Eq. (8). Suppose that $[A,B]\neq 0$ and we found such a state $|\phi\rangle=|\phi_{0}\rangle$, which is not an eigenstate of $A$ or of $B$, that $\langle\phi_{0}|[A,B]|\phi_{0}\rangle=0$. It appears that this property of the state $|\phi_{0}\rangle$ need not imply that $\delta_{\phi_{0}}A|\phi_{0}\rangle\perp\delta_{\phi_{0}}B|\phi_{0}\rangle$, i. e. that $\langle\phi_{0}|\delta_{\phi_{0}}A\,\delta_{\phi_{0}}B|\phi_{0}\rangle=0$. Indeed, from equations (Introduction) — (8) it follows that the result $\langle\phi_{0}|[A,B]|\phi_{0}\rangle=0$ is equivalent to the condition $\Im(\langle\phi_{0}|\delta_{\phi_{0}}A\;\delta_{\phi_{0}}B|\phi_{0}\rangle)=0$, but this does not mean that also $\Re(\langle\phi_{0}|\delta_{\phi_{0}}A\;\delta_{\phi_{0}}B|\phi_{0}\rangle)=0$ and it may happen that $\Re(\langle\phi_{0}|\delta_{\phi_{0}}A\;\delta_{\phi_{0}}B|\phi_{0}\rangle)\neq 0$ which implies that $\langle\phi_{0}|\delta_{\phi_{0}}A\;\delta_{\phi_{0}}B|\phi_{0}\rangle\neq 0$, and the number of such cases does not have to be small. On the other hand, from the same equations it follows that if $\langle\phi|\delta_{\phi}A\,\delta_{\phi}B|\phi\rangle=0$ then it must always be $\langle\phi|[A,B]|\phi\rangle=0$.
As an illustration of the case under consideration, let us analyze the following example: Assume that $A=\lambda_{3}$ and $B=\lambda_{4}$, where $\lambda_{3},\lambda_{4}$ are Gell–Mann matrices:
$\lambda_{3}=\left(\begin{array}[]{ccc}1&0&0\\ 0&-1&0\\ 0&0&0\\ \end{array}\right),\;\;\lambda_{4}=\left(\begin{array}[]{ccc}0&0&1\\ 0&0&0\\ 1&0&0\\ \end{array}\right),\;\;\lambda_{5}=\left(\begin{array}[]{ccc}0&0&i\\ 0&0&0\\ -i&0&0\\ \end{array}\right).$ (17)
Here $\lambda_{5}$ is also Gell–Mann matrix. They are self–adjoint and do not commute, $[\lambda_{3},\lambda_{4}]=-i\lambda_{5}\neq 0.$ For these matrices the inequality (5) takes the following form,
$\left\|\delta_{\phi}\lambda_{3}\,|\phi\rangle\right\|^{2}\,\cdot\,\left\|\delta_{\phi}\lambda_{4}|\phi\rangle\right\|^{2}\geq\left|\langle\phi|\delta_{\phi}\lambda_{3}\,\delta_{\phi}\lambda_{4}|\phi\rangle\right|^{2}.$ (18)
Let us choose for simplicity,
$|\phi\rangle=|\phi_{1}\rangle=\frac{1}{N}\,\left(\begin{array}[]{c}a\\ b\\ 0\\ \end{array}\right),$ (19)
where $N^{2}=|a|^{2}+|b|^{2}$ and $a,b\in\mathbb{C}$. Using (19) one gets
$\delta_{\phi_{1}}\lambda_{3}|\phi_{1}\rangle=\frac{1}{N^{3}}\,\left(\begin{array}[]{c}2|b|^{2}a\\ -2|a|^{2}b\\ 0\\ \end{array}\right)\neq 0,$ (20)
and
$\delta_{\phi_{1}}\lambda_{4}|\phi_{1}\rangle=\frac{1}{N}\,\left(\begin{array}[]{c}0\\ 0\\ a\\ \end{array}\right)\neq 0,$ (21)
(where $\delta_{\phi_{1}}\lambda_{k}|\phi_{1}\rangle=(\lambda_{k}-\langle\phi_{1}|\lambda_{k}|\phi_{1}\rangle)|\phi_{1}\rangle$ and $k=3,4$), which leads to the result $\langle\phi_{1}|\delta_{\phi_{1}}\lambda_{3}\;\delta_{\phi_{1}}\lambda_{4}|\phi_{1}\rangle=0$, and $(\Delta_{{\phi}_{1}}\lambda_{3})^{2}={\left\|\delta_{\phi_{1}}\lambda_{3}|\phi_{1}\rangle\right\|}^{2}\neq 0$, $(\Delta_{{\phi}_{1}}\lambda_{4})^{2}={\left\|\delta_{\phi_{1}}\lambda_{4}|\phi_{1}\rangle\right\|}^{2}\neq 0$. Hence one concludes that for $|\phi_{1}\rangle$ the inequality (18) and thus relations (3), (9), (10) take the following form,
$\Delta_{\phi_{1}}\lambda_{3}\,\cdot\,\Delta_{\phi_{1}}\lambda_{4}\geq 0.$ (22)
Note that this result holds for any $a\neq 0$ and $b\neq 0$ defining the vector $|\phi_{1}\rangle$. This shows that the number of such vectors may be very large.
Now let’s consider another example. Again, let $A=\lambda_{3}$ and $B=\lambda_{4}$. For these matrices the inequality (3) takes the following form,
$\Delta_{\phi}\lambda_{3}\,\cdot\,\Delta_{\phi}\lambda_{4}\geq\frac{1}{2}\,\left|\langle[\lambda_{3},\lambda_{4}]\rangle_{\phi}\right|\equiv\frac{1}{2}\,\left|\langle\lambda_{5}\rangle_{\phi}\right|,$ (23)
Assume that
$|\phi\rangle\;\;=\;\;|\phi_{2}\rangle=\frac{1}{\sqrt{3}}\,\left(\begin{array}[]{c}1\\ 1\\ 1\\ \end{array}\right),$ (24)
which leads to the result $\left|\langle\lambda_{5}\rangle_{\phi_{2}}\right|=0$, and hence one concludes that for $|\phi_{2}\rangle$ the inequality (23) takes the following form
$\Delta_{\phi_{2}}\lambda_{3}\,\cdot\,\Delta_{\phi_{2}}\lambda_{4}\geq 0,$ (25)
where $\Delta_{\phi_{2}}\lambda_{3}\neq 0$ and $\Delta_{\phi_{2}}\lambda_{4}\neq 0$.
Elementary calculations show that in this case
$\langle\phi_{2}|\delta_{\phi_{2}}A\;\delta_{\phi_{2}}B|\phi_{2}\rangle=\langle\phi_{2}|\delta_{\phi_{2}}\lambda_{3}\;\delta_{\phi_{2}}\lambda_{4}|\phi_{2}\rangle\equiv\langle\phi_{2}|\lambda_{3}\lambda_{4}|\phi_{2}\rangle=\frac{1}{3}.$ (26)
(It is because $\langle\phi_{2}|\lambda_{3}|\phi_{2}\rangle=0$). Finally we get
$\Delta_{\phi_{2}}\lambda_{3}\,\cdot\,\Delta_{\phi_{2}}\lambda_{4}\;\geq|\langle\phi_{2}|\delta_{\phi}\lambda_{3}\;\delta_{\phi}\lambda_{4}|\phi_{2}\rangle|=\frac{1}{3}\neq 0.$ (27)
Concluding: As can be seen from results (23) — (27), the existence of such a state $|\phi\rangle$ that $\langle[A,B]\rangle_{\phi}=0$ does not necessarily guarantee that the right–hand side of the more general inequalities (9), (10), (13) will also be equal to zero.
Discussion
Recently, uncertainty relations for sums of standard deviations or variances have been studied in the literature (see, e. g. (see, e. g. [10, 14, 12, 15, 16] and many other papers), which are considered to be more useful than the HR or Schrödinger uncertainty relations. These studies are motivated by the fact that the latter do not provide any bounds on the standard deviations of either the observables $A$ or $B$ when the system is in the state described by the eigenfunction of the observable $B$ (or $A$) which causes that the right-hand side of (3) inequality vanishes. Relations of this type are called "sum uncertainty relations". Let us now look at some of these "sum uncertainty relations" and see if they are sensitive to the effects described above. Their mathematical basis is the triangle inequality in the Hilbert state space: There is
$\left\|\,|\psi_{1}\rangle\right\|+\left\|\,|\psi_{2}\rangle\right\|\,\geq\,\left\|\,|\psi_{1}\rangle+|\psi_{2}\rangle\right\|,$ (28)
for each pair of vectors $|\psi_{1}\rangle,|\psi_{2}\rangle\in\cal{H}$. It is usually derived in Hilbert space using the Schwartz inequality (4) — see e. g. see, e. g., [6, 10, 11, 12]. If we now substitute $|\psi_{A}\rangle=\delta_{\phi}A|\phi\rangle$ and $|\psi_{B}\rangle=\delta_{\phi}B|\phi\rangle$ into (28) instead of $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, we get the "sum uncertainty relation" derived in [10],
$\left\|\delta_{\phi}A|\phi\rangle\right\|+\left\|\delta_{\phi}B\phi\rangle\right\|\,\geq\,\left\|\delta_{\phi}A|\phi\rangle+\delta_{\phi}B|\phi\rangle\right\|\equiv\left\|\delta_{\phi}(A+B)|\phi\rangle\right\|,$ (29)
that is,
$\Delta_{\phi}A+\Delta_{\phi}B\geq\Delta_{\phi}(A+B).$ (30)
When looking for inequalities of the "sum uncertainty relation" type, one can also use the triangle inequality of the second kind [13, 14]:
$\left\|\,|\psi_{1}\rangle\right\|^{2}+\left\|\,|\psi_{2}\rangle\right\|^{2}\,\geq\,\frac{1}{2}\left\|\,|\psi_{1}\rangle+|\psi_{2}\rangle\right\|^{2},$ (31)
which leads to the following sum uncertainty relation,
$(\Delta_{\phi}A)^{2}+(\Delta_{\phi}B)^{2}\geq\frac{1}{2}\left[\,\Delta_{\phi}(A+B)\right]^{2}.$ (32)
This type of generalization of the inequality (30) has been studied in many papers (see, e. g. [14, 15, 16, 17, 18, 19] and many others).
Now, let $|\phi\rangle=|\psi_{b}\rangle$, where $|\psi_{b}\rangle$ is and eigenvector of the observable $B$ for the eigenvalue $b$. Then, as it has been shown earlier, $\delta_{\psi_{b}}B|\psi_{b}\rangle=0$, and therefore $\Delta_{\psi_{b}}B=0$, and $\delta_{\psi_{b}}(A+B)|\psi_{b}\rangle=\delta_{\psi_{b}}A|\psi_{b}\rangle+\delta_{\psi_{b}}B|\psi_{b}\rangle\equiv\delta_{\psi_{b}}A|\psi_{b}\rangle$, and as a result we have $\Delta_{\psi_{b}}(A+B)\equiv\Delta_{\psi_{b}}A$. This means that even though the right–hand side of the inequalities (30) and (32) is non–zero, these inequalities become trivial: $\Delta_{\psi_{b}}A\geq\Delta_{\psi_{b}}A$ (or $(\Delta_{\psi_{b}}A)^{2}\geq\frac{1}{2}(\Delta_{\psi_{b}}A)^{2}$ respectively), and contains no information about $\Delta_{\psi_{b}}A$.
Let us now examine a case that is more interesting to us: Assume that $|\psi_{A}\rangle=\delta_{\phi}A|\phi\rangle\neq 0,\;|\psi_{B}\rangle=\delta_{\phi}B|\phi\rangle\neq 0$, and $\delta_{\phi}A|\phi\rangle\,\perp\,\delta_{\phi}B|\phi\rangle$. Analyzing the right hand sides of inequalities (31) and (32) with this assumption, one finds that $\left[\Delta_{\phi}(A+B)\right]^{2}\equiv\left\|\delta_{\phi}A|\phi\rangle+\delta_{\phi}B|\phi\rangle\right\|^{2}\equiv\left\|\delta_{\phi}A|\phi\rangle\right\|^{2}+\left\|\delta_{\phi}B\phi\rangle\right\|^{2}$. (It is because $\left\|\,|\psi_{A}\rangle+|\psi_{B}\rangle\right\|^{2}=\left\|\,|\psi_{A}\rangle\right\|^{2}+\left\|\,|\psi_{B}\rangle\right\|^{2}$ for any $|\psi_{A}\rangle\,\perp\,|\psi_{B}\rangle$). This means that finally, in the case under consideration, the inequality (32) takes the following form: $(\Delta_{\phi}A)^{2}+(\Delta_{\phi}B)^{2}\geq\frac{1}{2}\left[(\Delta_{\phi}A)^{2}+(\Delta_{\phi}B)^{2}\right]$ and we have no information about lower bounds for $(\Delta_{\phi}A)^{2}$ and $(\Delta_{\phi}B)^{2}$. Let us now analyze the inequality (30) using the assumption that $\delta_{\phi}A|\phi\rangle\,\perp\,\delta_{\phi}B|\phi\rangle$. We already know that then $\left[\Delta_{\phi}(A+B)\right]^{2}\equiv(\Delta_{\phi}A)^{2}+(\Delta_{\phi}B)^{2}$. Next, by multiplying the two sides of the inequality (30) by each other respectively, we get that $(\Delta_{\phi}A)^{2}+(\Delta_{\phi}B)^{2}+2\Delta_{\phi}A\,\Delta_{\phi}B\geq\left[\Delta_{\phi}(A+B)\right]^{2}\equiv(\Delta_{\phi}A)^{2}+(\Delta_{\phi}B)^{2}$. This inequality simplifies to $\Delta_{\phi}A\,\Delta_{\phi}B\geq 0$, which again gives us no information about the lower bounds on $\Delta_{\phi}A$ and $\Delta_{\phi}B$.
In summary, when we study two non–commuting observables, $A$ and $B$, and apply the "sum uncertainty relations" to find lower bounds on $\Delta_{\phi}A$ and $\Delta_{\phi}B$ in the situations considered above, the conclusions resulting from these relations are the same as the conclusions resulting from the HR uncertainty relations. The "sum uncertainty relations" seem to be more useful in the case where we investigate possible relations between lower bounds of the standard deviations, $(\Delta_{\phi}A_{1}),(\Delta_{\phi}A_{2}),\ldots,(\Delta_{\phi}A_{n})$ of a set of non–commuting observables $\left\{A_{j}\right\}_{j=1}^{n}$ with $n\geq 3$. In such cases, using the generalization of the triangle inequality to $n$ vectors (see, e. g., [11]),
$\sum_{j=1}^{n}\left\|\,|\psi_{j}\rangle\right\|\;\geq\;\left\|\sum_{j=1}^{n}|\psi_{j}\rangle\right\|,$ (33)
we obtain the following generalization of the inequality (29),
$\sum_{j=1}^{n}(\Delta_{\phi}A_{j})\geq\,\Delta_{\phi}(\sum_{j=1}^{n}A_{j}).$ (34)
and this type of "sum uncertainty relation" and its various generalizations analogous to (32) have recently been intensively studied in many papers (see, e. g. [10, 17, 18, 19] and many others).
Final remarks
From the analysis carried out in Sec. 2 it follows that the uncertainty relation in its the most general form (13) is in fact an upper bound on the modulus of the correlation function ${\cal C}_{\phi}(A,B)$. If $\Delta_{\phi}A>0$ and $\Delta_{\phi}B>0$ then the uncertainty relation (13) can be written in the equivalent form (14) as the quantum modification of the Pearson correlation coefficient $\mathfrak{r}_{\phi}(A,B)$. As it also has been shown there may exist large sets of states $S_{AB}=\{|\phi\rangle\in{\cal H}\,|\,\mathfrak{r}_{\phi}(A,B)=0\}\subset{\cal H}$, (where $\dim{\cal H}\geq 3$), of a quantum system, which are not eigenstates of any observable from non–commuting pairs $A$ and $B$, such that vectors $\delta_{\phi}A|\phi\rangle\neq 0$ and $\delta_{\phi}B|\phi\rangle\neq 0$ are orthogonal: $\delta_{\phi}A|\phi\rangle\,\perp\,\delta_{\phi}B|\phi\rangle$. These sets may be different for different pairs of such observables. It should be noted here that the set $S_{AB}$ is not identical to the set $S_{[A,B]}$ of all vectors for which the right side of the HR uncertainty relation (3) is equal to zero: $S_{[A,B]}=\{|\phi\rangle\in{\cal H}\,|\,\langle\phi|[A,B]|\phi\rangle=0\,{\rm and}\,[A,B]\neq 0\}$. There is $S_{A,B}\subset S_{[A,B]}$ and $S_{[A,B]}\setminus S_{A,B}\neq\emptyset$ (see Sec. 2). On vectors belonging to the set $S_{AB}$ the right–hand sides of the inequalities (3), (5) — (10) reach their absolute minimum, i.e. they have the value zero. So, there is no a lower bound for the product $\Delta_{\phi}A\cdot\Delta_{\phi}B$. Strictly speaking, this lower bound is zero. In other words, quantum theory (to be more precisely: basic assumptions (postulates) of the quantum theory) allows $\Delta_{\phi}A$ and $\Delta_{\phi}B$ both to be as small as possible and also it allows that in this case the magnitude of $\Delta_{\phi}A$ does not affect the magnitude of $\Delta_{\phi}B$ and vice versa. Non–commuting observables $A$ and $B$ are uncorrelated if the system is in a state $|\phi\rangle\in S_{AB}$ and then they behave as if they were independent: In such cases, observable $A$ does not disturb observable $B$ (and vice versa) and does not generate any additional fluctuations of $B$. This raises the questions: Can this property be used technically in any way, and if so, how? What are properties of the system prepared in such a state?